Could someone please help me with setting this problem up and then how to begin to solve it? I have read my algebra book over and over and still can't get it. Thanks so much. A coffee merchant wants to make 25lb of a blend of coffee to sell for $4.25 per pound. The blend is made by using a $6 grade and a $3.50 grade of coffee. How many pounds of each grade of coffee should be used? Thanks for your time
Answer by stanbon(75887)

A coffee merchant wants to make 25lb of a blend of coffee to sell for $4.25 per pound. The blend is made by using a $6 grade and a $3.50 grade of coffee. How many pounds of each grade of coffee should be used?
----------------
$6 grade coffee DATA:
Amount = x lbs. ; value = 6x dollars
------------------
$3.50 grade cofee DATA:
Amount = "25-x" lbs; value = 3.5(25-x)= 87.5-3.5x dollars
--------------------
$4.25 mixture DATA:
Amount = 25 lbs ; value = 4.25*25 = $106.25
--------------------
EQUATION:
value + value = 106.25 dollars
6x + 87.5-3.5x = 106.25
2.5x = 18.75
x = 7.5 lbs ( amount of $6 coffee in the blend)
25-x = 17.5 ( amount of #3.50 coffee in the blend)
====================
Cheers,
Stan H.